# Aspects of the geometry and topology of expanding horizons

### v1

@@ Abstract.
The aim of this paper is to extend some basic results about marginally outer trapped surfaces to the context of surfaces having general null expansion. Motivated in part by recent work of Chai-Wan, we introduce the notion of $\mathfrak{g}$-stability for a general closed hypersurface $\Sigma$ in an ambient initial data set and prove that, under natural energy conditions, $\Sigma$ has positive Yamabe type, that is, $\Sigma$ admits a metric of positive scalar curvature, provided $\Sigma$ is $\mathfrak{g}$-stable. Similar results are obtained when $\Sigma$ is embedded in a spacelike, or null, hypersurface of a spacetime satisfying the dominant energy condition. Conditions implying $\mathfrak{g}$-stability are also discussed. Finally, we obtain a spacetime positive mass theorem for initial data sets with compact boundary $\Sigma$ of positive null expansion, assuming that the dominant energy condition is sufficiently strict near $\Sigma$. This extends recent results of Galloway-Lee and Lee-Lesourd-Unger.
@@ 1. Introduction
Marginally outer trapped surfaces are objects of considerable interest at the interface of spacetime geometry and the physics of black holes. The notion of a marginally outer trapped surface (or MOTS for short) was introduced early on in the development of the theory of black holes, in connection with gravitational collapse; see e.g [16]. MOTSs appeared in a more purely mathematical context in the work of R. Schoen and S.-T. Yau [22] concerning the existence of solutions of Jang’s equation, in connection with their proof of the positivity of energy. MOTSs may be viewed as spacetime analogues of minimal surfaces in Riemannian geometry and, despite the absence of a variational characterization like that for minimal surfaces, satisfy a number of analogous properties.
MOTSs arise in various situations. For example, cross-sections of the event horizon in stationary black hole spacetimes (such as Schwarzschild and Kerr) are MOTSs. This can be roughly understood in terms of Hawking’s area theorem. For dynamically evolving black hole spacetimes, the null geodesic generators of the event horizon have nonnegative expansion towards the future. However, in the steady state (stationary) limit, this expansion goes to zero. In dynamical black hole spacetimes (such as the Vaidya spacetime with null dust source), MOTSs typically occur inside the event horizon (as the boundary of the trapped region within a spacelike slice). In fact, there are general results showing that under appropriate conditions, MOTSs cannot occur outside the event horizon. (See Section 2 for a brief discussion.)
@@
Our first three results extend results on the topology of MOTSs [11, 12, 14] to surfaces of general nonnegative null expansion. Definitions of various terms will be given in the next section. The first result is a pure initial data result.
Theorem 1.1.
Let $(M,g,K)$ be an initial data set and $\Sigma$ be a closed hypersurface in $M$. If $\Sigma$ is a $\mathfrak{g}$-stable surface in $(M,g,K)$ with null expansion $\theta^{+}=h\in C^{\infty}(\Sigma)$, $h\geq 0$ and either
  1. (i)
    $\mu-|J|\geq 0$ and $\tau\leq 0$ along $\Sigma$, $h\not\equiv 0$, or
  2. (ii)
then $\Sigma$ admits a metric of positive scalar curvature.
As is well-known there are many restrictions on the topology of manifolds that admit a metric of positive scalar curvature (see e.g. [8] for a recent reference).
We note that a result related to Theorem 1.1 was obtained in [7] using a slightly different version of stability, which is well-adapted to constructing surfaces of prescribed null expansion.
The next two results are spacetime results, and only require the (spacetime) dominant energy condition; no strictness is needed. They assume that $\Sigma$ is a surface in a spacelike, or null, hypersurface, respectively. Physically, we may think of $\Sigma$ as arising from the intersection of the event horizon (assumed to be sufficiently smooth) with these hypersurfaces.
Theorem 1.2.
Let $(\bar{M},\bar{g})$ be a spacetime satisfying the dominant energy condition and $M$ be a spacelike hypersurface in $(\bar{M},\bar{g})$. If $\Sigma$ is a $\mathfrak{g}$-locally weakly outermost surface in $M$ with constant null expansion $\theta^{+}=\theta_{0}>0$, then $\Sigma$ admits a metric of positive scalar curvature.
The next result assumes a natural condition on the inner null expansion of $\Sigma$.
Theorem 1.3.
Our final theorem pertains to the spacetime positive mass theorem [9]. Recent results have extended this theorem to initial data sets with boundary [13, 18], where the components are assumed to be weakly outer trapped ($\theta^{+}\leq 0$), or weakly inner untrapped ($\theta^{-}\geq 0$). Here we show that the weakly outer trapped condition can be relaxed a bit, by requiring the dominant energy condition (for initial data sets) to be sufficiently strict near $\Sigma$.
Theorem 1.4.
Proofs are presented in Section 3. In the next section, we introduce concepts pertinent to surfaces of general null expansion.
@@ Acknowledgements.
The work of GJG was partially supported by the Simons Foundation, under Award No. 850541. The work of AM was partially supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq, Brazil (Grants 305710/2020-6, 405468/2021-0, and 309867/2023-1) and the Fundação de Amparo à Pesquisa do Estado de Alagoas - Fapeal, Brazil (Process No. E:60030.0000002254/2022).
2. Preliminaries
@@
The local energy density $\mu$ and the local current density $J$ of an initial data set $(M,g,K)$ are given by
@@
where $S$ is the scalar curvature of $(M,g)$. When $(M,g,K)$ is a spacetime initial data set, these quantities are given by $\mu=G(u,u)$ and $J=G(u,\cdot)$, where $G$ is the Einstein tensor $G={\rm Ric}_{\bar{M}}-\frac{1}{2}R_{\bar{M}}\bar{g}$. Moreover, the inequality,
@@
Though not strictly necessary, for the purpose of introducing certain concepts, we shall assume $(M,g,K)$ is an initial data set in a spacetime $(\bar{M},\bar{g})$. (In fact, this can always be arranged, cf. [2].) Then the second fundamental form $K$ of $M$ is given by,
@@
We now introduce a notion of stability for surfaces $\Sigma$ of prescribed null expansion $\theta^{+}=h$, which generalizes in a straightforward way the usual notion of stability of marginally outer trapped surfaces ($\theta^{+}=0$). (As noted in the introduction, a slightly different notion of stability for surfaces of prescribed null expansion was considered in [7]; see also [6].)
@@
$\Psi^{u}:[0,1]\times\Sigma\to M,\quad\Psi^{u}(t,p)=\exp_{p}(tu(p)\nu(p)),$ (2.2)
@@
where $L:C^{\infty}(\Sigma)\to C^{\infty}(\Sigma)$ is the elliptic operator
@@
where $h=\theta_{u}^{+}(0)$ is the null expansion of $\Sigma_{0}^{u}=\Sigma$,
@@
$\gamma=\langle\,,\,\rangle$ is the induced metric on $\Sigma$, $R_{\gamma}$ is the scalar curvature of $\gamma$, and $X$ is the vector field tangent to $\Sigma$ that is dual to the 1-form $K(\nu,\cdot)|_{\Sigma}$. Moreover, there is a real number $\lambda$, called the principal eigenvalue of $L$, satisfying $L\phi=\lambda\phi$ for some positive eigenfunction $\phi\in C^{\infty}(\Sigma)$, such that $\lambda\leq\mbox{Re}(\mu)$ for any other eigenvalue $\mu$ of $L$. Also, the eigenspace of $L$ associated with $\lambda$ has dimension 1. By arguments essentially as in [1] (see also [17, Theorem A.10]), one has that $\Sigma$ is $\mathfrak{g}$-stable if, and only if, $\lambda\geq 0$.
@@
$\theta^{+}_{u}(t)(p_{t,u})<h(p)\quad\text{for all}\quad p\in\Sigma,$ (2.4)
where $p_{t,u}=\Psi^{u}(t,p)$. If $\Sigma$ is $\mathfrak{g}$-locally weakly outermost then it is necessarily $\mathfrak{g}$-stable. In fact, if $\Sigma$ is not $\mathfrak{g}$-stable, then $\lambda<0$. Let $\phi>0$ be a principal eigenfunction of $L$. Without loss of generality, we may assume that $\|\phi\|_{C^{0}}\leq\epsilon_{0}$. Therefore,
@@
We mention a simple condition that implies the $\mathfrak{g}$-locally weakly outermost condition.
@@
Let $\Sigma$ be a closed embedded hypersurface in an initial data set $(M,g,K)$ with null expansion $\theta^{+}=h$. Suppose there exists a variation $\Psi^{\hat{u}}$ of $\Sigma$, $\|\hat{u}\|_{C^{0}}\leq\hat{\epsilon}_{0}$, such that $\theta^{+}_{\hat{u}}(\hat{t})\geq h$ pointwise for all $\hat{t}\in[0,1]$. Then $\Sigma$ is $\mathfrak{g}$-locally weakly outermost.
@@
Let $\Psi^{u}$, $u>0$, $\|u\|_{C^{0}}\leq\epsilon_{0}$, be any variation of $\Sigma$. Choose $\epsilon_{0}$ sufficiently small so that $\Psi^{u}([0,1]\times\Sigma)\subset\Psi^{\hat{u}}([0,1]\times\Sigma)$. Suppose for some $t\in[0,1]$, $\theta^{+}_{u}(t)<h$ pointwise. By the compactness of $\Sigma_{t}^{u}$, there exists $\hat{t}$ such that $\Sigma_{t}^{u}$ lies to the inside of $\Sigma_{\hat{t}}^{\hat{u}}$ and so that they meet tangentially at some point $q=\Psi^{u}(t,p)$. Restricting the size of $\Sigma$ to a small neighborhood of $p$, we may assume there exists a constant $a$ such that $\theta_{u}^{+}(t)<a<\theta^{+}_{\hat{u}}(\hat{t})$. But then the maximum principle for null expansion ([2, Prop. 2.4], [4, Prop. 3.1]) would require $\theta^{+}_{u}(t)=\hat{\theta}^{+}_{\hat{u}}(\hat{t})$, a contradiction.∎
For example, the CMC spheres $r=r_{0}$ between the horizon and photon sphere ($2m<r_{0}<3m$) in the totally geodesic ($K=0$) time slice $t=t_{0}$ of Schwarzschild spacetime satisfy the conditions of the proposition and hence are $\mathfrak{g}$-locally weakly outermost. At the end of this section, we include further comments about the $\mathfrak{g}$-locally weakly outermost condition.
We now extend the concept of $\mathfrak{g}$-stability for surfaces in initial data sets (e.g. in spacelike hypersurfaces) to surfaces in null hypersurfaces. Let $\mathcal{N}$ be a null hypersurface in a spacetime $(\bar{M},\bar{g})$ and $\Sigma$ be a closed hypersurface in $\mathcal{N}$ that is spacelike in $(\bar{M},\bar{g})$. Fix a future-pointing null vector field $\ell^{-}$ on $\Sigma$ that is orthogonal to $\Sigma$ and tangent to $\mathcal{N}$.
@@
is well defined. Now, extend $\ell^{-}$ to a neighborhood of $\Sigma$ in $\mathcal{N}$ by
@@
and define $\ell^{+}$ as the future-pointing null vector field that is normal to $\mathcal{N}$ and satisfies $\langle\ell^{+},\ell^{-}\rangle=-2$. In this case, the null expansion scalars $\theta^{+}$, $\theta^{-}$ of $\Sigma$ in $\mathcal{N}$ are defined by
@@
$F^{u}:[0,1]\times\Sigma\to\mathcal{N},\quad F^{u}(t,p)=\exp_{p}(-tu(p)\ell^{-}(p))$
is also well defined for $u\in C^{\infty}(\Sigma)$ positive with $\|u\|_{C^{0}}\leq\epsilon_{0}$. Denote $F^{u}(t,\Sigma)$ by $\Sigma_{t}^{u}$ and the null expansion scalars of $\Sigma_{t}^{u}$ in $\mathcal{N}$ by $\theta_{u}^{\pm}(t)=\operatorname{div}_{\Sigma_{t}^{u}}\ell^{\pm}$. In this case, from well-known formulas ([1, 3]), we have,
@@
where $L_{-}:C^{\infty}(\Sigma)\to C^{\infty}(\Sigma)$ is the elliptic operator
@@
Again, as in the initial data case, one has that $\Sigma$ is $\mathfrak{g}$-stable if, and only if, $\lambda_{1}(L_{-})\geq 0$, where $\lambda_{1}(L_{-})$ is the principal eigenvalue of $L_{-}$. A completely analogous definition of $\mathfrak{g}$-locally weakly outermost holds in this null hypersurface case (just replace $\Psi^{u}$ by $F^{u}$), and again $\mathfrak{g}$-locally weakly outermost implies $\mathfrak{g}$-stability.
@@
Let $\Sigma$ be a closed embedded hypersurface in a null hypersurface $\mathcal{N}$. Suppose there exists a variation $F^{\hat{u}}$ of $\Sigma$, $\|\hat{u}\|_{C^{0}}\leq\hat{\epsilon}_{0}$, such that $\theta^{+}_{\hat{u}}(\hat{t})\geq h$ pointwise for all $\hat{t}\in[0,1]$. Then $\Sigma$ is $\mathfrak{g}$-locally weakly outermost.
The proof is essentially the same as in Proposition 2.1, except a variation of the maximum principle referenced there is needed, one that applies to null, rather than spacelike, hypersurfaces. That the maximum principle for null expansion holds in this case follows from [20, Theorem 2], which is a consequence of the maximum principle for null hypersurfaces [10].
Vaidya spacetime, which may be viewed as a dynamical version of Schwarzschild spacetime, provides a nice illustration of Proposition 2.2. Using formulas in [5] for the Vaidya spacetime in ingoing Eddington-Finkelstein coordinates $(v,r,\theta,\phi)$, together with the approximate location of the event horizon $H$ (as analyzed in e.g. [21]), it can be shown for suitable mass functions that spherically symmetric cross-sections of $H$ (which will have constant positive outward null expansion) satisfy the conditions of Proposition 2.2, and hence are $\mathfrak{g}$-locally weakly outermost.
The following will be important for some of our results. It was proved in [11] (based on the main argument in [14]).
Lemma 2.3.
@@
where $X$ is a tangent vector field on $\Sigma$ and $\mathcal{Q}=\frac{1}{2}R_{\gamma}-\mathcal{P}$ for some function $\mathcal{P}\geq 0$. If $\lambda_{1}(L)\geq 0$, then $\Sigma$ admits a metric of positive scalar curvature, unless $\lambda_{1}(L)=0$, $\mathcal{P}\equiv 0$, and $(\Sigma,\gamma)$ is Ricci flat.
Further comments on the $\mathfrak{g}$-locally weakly outermost condition. It is a basic result in the theory of black holes that for suitably defined black hole spacetimes satisfying the null energy condition (NEC), ${\rm Ric}(X,X)\geq 0$ for all null vectors $X$, outer trapped surfaces cannot exist outside the event horizon. See, for example, [23, Prop. 12.2.2], which, under natural circumstances, extends to outer trapped (not just trapped) surfaces. This provides a certain rationale for assuming that a MOTS is weakly outermost, and hence stable. The proof is based on the following. Suppose $\Sigma^{\prime}$ is an outer trapped surface outside the event horizon. Then one can construct an outgoing future complete null normal geodesic $\eta$ to $\Sigma^{\prime}$ without null focal points. But the assumption $\theta^{+}<0$, together with the NEC and the future completeness of $\eta$, imply that $\eta$ must necessarily contain a focal point.
The result above can be modified in a way that takes into account the presence of matter. Suppose, instead of the NEC, the following holds: along each future complete outward null normal geodesic $s\to\eta(s)$ to $\Sigma^{\prime}$, with $\eta^{\prime}(0)=\ell^{+}(p)$, $p=\eta(0)\in\Sigma^{\prime}$,

### v2

@@ Abstract.
The aim of this paper is to extend some basic results about marginally outer trapped surfaces to the context of surfaces having general null expansion. Motivated in part by recent work of Chai-Wan, we introduce the notion of $\mathfrak{g}$-stability for a general closed hypersurface $\Sigma$ in an ambient initial data set and prove that, under natural energy conditions, $\Sigma$ has positive Yamabe type, that is, $\Sigma$ admits a metric of positive scalar curvature, provided $\Sigma$ is $\mathfrak{g}$-stable. A similar result is obtained when $\Sigma$ is embedded in a null hypersurface of a spacetime satisfying the dominant energy condition. Area bounds under similar conditions are obtained in the case where $\Sigma$ is $2$-dimensional. Conditions implying $\mathfrak{g}$-stability are also discussed. Finally, we obtain a spacetime positive mass theorem for initial data sets with compact boundary $\Sigma$ of positive null expansion, assuming that the dominant energy condition is sufficiently strict near $\Sigma$. This extends recent results of Galloway-Lee and Lee-Lesourd-Unger.
@@ 1. Introduction
Marginally outer trapped surfaces are objects of considerable interest at the interface of spacetime geometry and the physics of black holes. The notion of a marginally outer trapped surface (or MOTS for short) was introduced early on in the development of the theory of black holes, in connection with gravitational collapse; see e.g. [18]. MOTSs appeared in a more purely mathematical context in the work of R. Schoen and S.-T. Yau [25] concerning the existence of solutions of Jang’s equation, in connection with their proof of the positivity of energy. MOTSs may be viewed as spacetime analogues of minimal surfaces in Riemannian geometry and, despite the absence of a variational characterization like that for minimal surfaces, satisfy a number of analogous properties.
MOTSs arise in various situations. For example, under suitable conditions, cross-sections of the event horizon in stationary black hole spacetimes (such as Schwarzschild and Kerr) are MOTSs. This can be roughly understood in terms of Hawking’s area theorem. For dynamically evolving black hole spacetimes, the null geodesic generators of the event horizon have nonnegative expansion towards the future. However, in the steady state (stationary) limit, this expansion goes to zero. In dynamical black hole spacetimes (such as the Vaidya spacetime with null dust source), MOTSs typically occur inside the event horizon (as the boundary of the trapped region within a spacelike slice). In fact, there are general results showing that under appropriate conditions, MOTSs cannot occur outside the event horizon. (See Section 2.3 for a brief discussion.)
@@
In the next section, we present in some detail the key notions of stability for surfaces in expanding horizons, which we use in order to prove some results about the topology and area of these surfaces. These results are presented in Sections 3.1 and 3.2. In Section 3.3, we consider a version of the spacetime positive mass theorem that applies to expanding horizons.
@@ Acknowledgements.
The authors would like to thank the referee for many valuable comments on the first version of the paper. The work of the first named author was partially supported by the Simons Foundation, under Award No. 850541. The work of the second author was partially supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq, Brazil (Grants 309867/2023-1 and 405468/2021-0) and the Fundação de Amparo à Pesquisa do Estado de Alagoas - FAPEAL, Brazil (Process E:60030.0000002254/2022).
2. Surfaces of general null expansion: stability properties
@@
The local energy density $\mu$ and the local current density $J$ of an initial data set $(M,g,K)$ are given by,
@@
where $S$ is the scalar curvature of $(M,g)$. When $(M,g,K)$ is a spacetime initial data set, these quantities are given by $\mu=G(u,u)$ and $J=G(u,\cdot)$, where $G$ is the Einstein tensor $G={\rm Ric}_{\bar{M}}-\frac{1}{2}R_{\bar{M}}\bar{g}$ of $(\bar{M},\bar{g})$. Moreover, the inequality,
@@
Though not strictly necessary, for the purpose of introducing certain concepts, we shall assume that $(M,g,K)$ is an initial data set in a spacetime $(\bar{M},\bar{g})$. (In fact, this can always be arranged, cf. [2].) Then the second fundamental form $K$ of $M$ is given by,
@@
2.1. Stability with respect to initial data sets
We now introduce a notion of stability for surfaces $\Sigma$ of general null expansion $\theta^{+}=h$, which extends in a straightforward way the usual notion of stability of marginally outer trapped surfaces ($\theta^{+}=0$) in spacelike hypersurfaces. (As noted in the introduction, a slightly different notion of stability for surfaces of prescribed null expansion was considered in [7]; see also [6].) In order to discuss the stability properties of such surfaces (not necessarily MOTSs), we introduce a notation that allows one to compare pointwise the null expansions of two different surfaces.
@@ 2.1. Stability with respect to initial data sets
$\Psi^{u}:[0,1]\times\Sigma\to M,\quad\Psi^{u}(t,p)=\exp_{p}(tu(p)\nu(p)),$
@@
Heuristically speaking, $\Sigma$ is $\mathfrak{g}$-stable if there is an outward geodesic normal variation such that the null expansion is “infinitesimally nondecreasing”. For MOTS, stability has essentially the same meaning, and is described by saying that a stable MOTS is “infinitesimally outermost”.
@@
where $L:C^{\infty}(\Sigma)\to C^{\infty}(\Sigma)$ is the elliptic operator,
@@
where $h=\theta_{u}^{+}(0)$ is the null expansion of $\Sigma_{0}^{u}=\Sigma$, $\tau={\rm tr}K$ is the mean curvature of $M$,
@@
$\gamma=\langle\,,\,\rangle$ is the induced metric on $\Sigma$, $R_{\gamma}$ is the scalar curvature of $\gamma$, and $X$ is the vector field tangent to $\Sigma$ that is dual to the 1-form $K(\nu,\cdot)|_{\Sigma}$. Moreover, there is a real number $\lambda$, called the principal eigenvalue of $L$, satisfying $L\phi=\lambda\phi$ for some positive eigenfunction $\phi\in C^{\infty}(\Sigma)$, such that $\lambda\leq\mbox{Re}(\mu)$ for any other eigenvalue $\mu$ of $L$. Also, the eigenspace of $L$ associated with $\lambda$ has dimension 1. By arguments essentially as in [1] (see also[20, Theorem A.10]), one has that $\Sigma$ is $\mathfrak{g}$-stable if, and only if, $\lambda\geq 0$.
@@
$\theta^{+}_{u}(t)(p_{t,u})<h(p)\quad\text{for all}\quad p\in\Sigma,$
where $p_{t,u}=\Psi^{u}(t,p)$. If $\Sigma$ is $\mathfrak{g}$-locally weakly outermost, then it is necessarily $\mathfrak{g}$-stable. In fact, if $\Sigma$ is not $\mathfrak{g}$-stable, then $\lambda<0$. Let $\phi>0$ be a principal eigenfunction of $L$. Without loss of generality, we may assume that $\|\phi\|_{C^{0}}\leq\epsilon_{0}$. Therefore,
@@
We mention a simple criterion for the $\mathfrak{g}$-locally weakly outermost condition.
@@
Let $\Sigma$ be a closed embedded hypersurface in an initial data set $(M,g,K)$ with null expansion $\theta^{+}=h$. Suppose there exists a variation $\Psi^{\hat{u}}$ of $\Sigma$, $\hat{u}>0$, $\|\hat{u}\|_{C^{0}}\leq\hat{\epsilon}_{0}$, such that $\theta^{+}_{\hat{u}}(\hat{t})\geq h$ pointwise for all $\hat{t}\in[0,1]$. Then $\Sigma$ is $\mathfrak{g}$-locally weakly outermost.
@@
Let $\Psi^{u}$, $u>0$, $\|u\|_{C^{0}}\leq\epsilon_{0}$, be any variation of $\Sigma$. Choose $\epsilon_{0}$ sufficiently small so that $\Psi^{u}([0,1]\times\Sigma)\subset\Psi^{\hat{u}}([0,1]\times\Sigma)$. Suppose, by contradiction, that $\theta^{+}_{u}(t)<h$ pointwise, for some $t\in[0,1]$. By the compactness of $\Sigma_{t}^{u}$, there exists $\hat{t}$ such that $\Sigma_{t}^{u}$ lies to the inside of $\Sigma_{\hat{t}}^{\hat{u}}$ and so that they meet tangentially at some point $q=\Psi^{u}(t,p)$. Restricting the size of $\Sigma$ to a small neighborhood of $p$, we may assume there exists a constant $a$ such that $\theta_{u}^{+}(t)<a<\theta^{+}_{\hat{u}}(\hat{t})$. But then the maximum principle for null expansion ([2, Prop. 2.4],[4, Prop. 3.1]) would require $\theta^{+}_{u}(t)=\hat{\theta}^{+}_{\hat{u}}(\hat{t})$, which is a contradiction. ∎
We mention several examples. Consider the CMC spheres $r=r_{0}$ between the horizon and photon sphere ($2m<r_{0}<3m$) in the totally geodesic ($K=0$) time slice $t=t_{0}$ of Schwarzschild spacetime. The mean curvature (and hence the null expansion) of these spheres increases as $r$ increases from $2m$ to $3m$. It then follows from the proposition that each such sphere is $\mathfrak{g}$-locally weakly outermost, and hence stable. The fact that they are spheres is consistent with Theorem 3.1.
Now consider the manifold $M=\mathbb{R}\times V$, with metric $g=dt^{2}+\cosh^{2}t\,g_{\Sigma}$, where $(\Sigma,g_{\Sigma})$ is a higher genus surface of constant Gaussian curvature $-1$. Then $M$ has constant curvature $-1$. Consider the initial data set $(M,g,K=g)$, which one easily verifies is a vacuum initial data set; in fact, $\mu=0$ and $J=0$. One further checks that the null expansion $h_{t}$ of the surfaces $\Sigma_{t}=\{t\}\times\Sigma$ increases from 00 to $4$ as $t$ increases from $\infty$- ∞ to +$\infty$+ ∞. It again follows from the proposition that each $\Sigma_{t}$ is $\mathfrak{g}$-locally weakly outermost, and hence stable. That the $\Sigma_{t}$’s do not admit metrics of positive scalar curvature is consistent with Theorem 3.1. In this example, $\tau=3>\frac{1}{2}h_{t}$, and so illustrates the relevance of condition (i) in Theorem 3.1. If we take $(M,g,K=ag)$, for any $a>1$, we have $\mu=3(a^{2}-1)$ and $J=0$. Now, $h_{t}$ increases from $-2+2a$ to $2+2a$ as $t$ increases from $\infty$- ∞ to +$\infty$+ ∞. In this case, $h_{t}\tau\geq 6a(a-1)>3(a^{2}-1)$. This example illustrates the relevance of condition (ii) in Theorem 3.1.
2.2. Stability with respect to null hypersurfaces
We now extend the concept of $\mathfrak{g}$-stability for surfaces in initial data sets (e.g. in spacelike hypersurfaces) to surfaces in null hypersurfaces. Let $\mathcal{N}$ be a null hypersurface in a spacetime $(\bar{M},\bar{g})$ and $\Sigma$ be a closed hypersurface in $\mathcal{N}$ that is spacelike in $(\bar{M},\bar{g})$. Fix a future-pointing null vector field $\ell^{-}$ along $\Sigma$ that is orthogonal to $\Sigma$ and tangent to $\mathcal{N}$.
@@ 2.2. Stability with respect to null hypersurfaces
is well defined. Now, extend $\ell^{-}$ to a neighborhood of $\Sigma$ in $\mathcal{N}$ by,
@@
and define $\ell^{+}$ as the future-pointing null vector field along $\Sigma$ that is normal to $\Sigma$ and satisfies $\langle\ell^{+},\ell^{-}\rangle=-1$. In this case, the null expansion scalars $\theta^{+}$, $\theta^{-}$ of $\Sigma$ in $\mathcal{N}$ are defined by,
@@
$F^{u}:[0,1]\times\Sigma\to\mathcal{N},\quad F^{u}(t,p)=\exp_{p}(-tu(p)\ell^{-}(p)),$
is also well defined for $u\in C^{\infty}(\Sigma)$ positive with $\|u\|_{C^{0}}\leq\epsilon_{0}$. Extend $\ell^{+}$ to be the future-pointing null vector field that is normal to $\Sigma_{t}^{u}:=F^{u}(t,\Sigma)$ for each $t$ and satisfies $\langle\ell^{+},\ell^{-}\rangle=-1$. Denote the null expansion scalars of $\Sigma_{t}^{u}$ in $\mathcal{N}$ by $\theta_{u}^{\pm}(t)=\operatorname{div}_{\Sigma_{t}^{u}}\ell^{\pm}$. In this case, from well-known formulas ([1, 3]), we have,
@@
where $L_{-}:C^{\infty}(\Sigma)\to C^{\infty}(\Sigma)$ is the elliptic operator,
@@
Similarly to the initial data case, we make the following definition.
Definition 2.3.
@@
[…]
@@
Let $\Sigma$ be a closed embedded hypersurface in a null hypersurface $\mathcal{N}$. Suppose there exists a variation $F^{\hat{u}}$ of $\Sigma$, $\hat{u}>0$, $\|\hat{u}\|_{C^{0}}\leq\hat{\epsilon}_{0}$, such that $\theta^{+}_{\hat{u}}(\hat{t})\geq h$ pointwise for all $\hat{t}\in[0,1]$. Then $\Sigma$ is $\mathfrak{g}$-locally weakly outermost.
The proof is essentially the same as in Proposition 2.2, except that a variation of the maximum principle referenced there is needed, one that applies to null, rather than spacelike, hypersurfaces. That the maximum principle for null expansion holds in this case follows from [23, Theorem 2], which is a consequence of the maximum principle for null hypersurfaces [11].
Vaidya spacetime, which may be viewed as a dynamical version of Schwarzschild spacetime, provides a nice illustration of Proposition 2.4. The metric in ingoing Eddington-Finkelstein coordinates $(v,r,\theta,\phi)$ is given by the line element,
$ds^{2}=-\left(1-\frac{2M}{r}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2},$
where $M=M(v)$. The dominant energy condition is satisfied if $M(v)$ is nondecreasing. Consider a null hypersurface $\mathcal{N}:v=v_{0}$. The $2$-spheres $\Sigma_{r}$ in $\mathcal{N}$ have null expansion
$\theta^{+}_{r}=\frac{r-2M}{r^{2}}$
with respect to the future directed outward null normal,
$\ell^{+}=\frac{\partial}{\partial v}+\frac{1}{2}\left(1-\frac{2M}{r}\right)\frac{\partial}{\partial r},$
where $M=M(v_{0})$ (see e.g. [5]). The null expansion $\theta^{+}_{r}$ increases as $r$ increases from $2M$ to $4M$. Using the approximate location of the event horizon $\mathcal{H}$ (as analyzed in e.g. [24]), it can be shown that, for suitable functions $M=M(v)$, the spherically symmetric intersection of $\mathcal{H}$ and $\mathcal{N}$ will lie well within this range. It then follows from Proposition 2.4 that this intersection is $\mathfrak{g}$-locally weakly outermost, and hence stable.
The following will be important for the proofs of several of our results. It follows from a straightforward generalization of arguments in [16]; see [12, Sect. 2] and [14, Sect. 2].
Theorem 2.5.
@@
where $\mathcal{Q}\in C^{\infty}(\Sigma)$ and $X$ is a smooth tangent vector field on $\Sigma$. Suppose the principal eigenvalue of $L$ is nonnegative, $\lambda_{1}(L)\geq 0$. Then the following hold:
  1. (i)